I got the answer correct but I don't know why. Here's the question:The table below shows the total value (in dollars) that Andrew gets from playing 9-hole rounds of golf. Rounds of Golf per Mo
Answer #1

Answer : The answer is option D.

In case of 0 rounds, the cost = Price of golf * Number of rounds = 13 * 0 = $0. So, in this case profit = Total value - Cost = 0 - 0 = $0.

In case of 1 rounds, the cost = Price of golf * Number of rounds = 13 * 1 = $13. So, in this case profit = Total value - Cost = 40 - 13 = $27.

In case of 2 rounds, the cost = Price of golf * Number of rounds = 13 * 2 = $26. So, in this case profit = Total value - Cost = 70 - 26 = $44.

In case of 3 rounds, the cost = Price of golf * Number of rounds = 13 * 3 = $39. So, in this case profit = Total value - Cost = 92 - 39 = $53.

In case of 4 rounds, the cost = Price of golf * Number of rounds = 13 * 4 = $52. So, in this case profit = Total value - Cost = 108 - 52 = $56

Here we can see that Andrew gets more profit by playing 4 rounds per month. Hence option D is the correct answer.